# Continuous-wave, high-resolution, ultra-broadband mid-infrared nonlinear spectroscopy with tunable plasmonic nanocavities

### v1

@@ Results
Several open questions remain. First, the SFG spectrum from BPhT presents significant deviations from that predicted by DFT calculations. Could this be related to the embedding of the molecules in a plasmonic nanogap? Is the discrepancy a signature of vacuum electromagnetic field modification [50], or rather linked to chemical interaction altering the geometry or electronic states of the molecules? Or is it related to other contributions not yet accounted for, such as from charge-transfer resonances? Unfortunately, we could not find broadband SFG spectra of BPhT from flat gold films in the literature to serve as a comparison. We attempted the DFT simulation of a more realistic geometry with molecules sandwiched between two gold planes using the model of Ref. [53], yet this did not yield better agreement with the data. Such complex metal-molecule interfaces are still a challenging system to accurately model with DFT.
@@
Second, the strength of the non-resonant varies markedly compared to the resonant vibrational response, when changing the molecular species. This observation challenges the assumption that the non-resonant is an intrinsic property of the gold surface [54]. We speculate that the existence of charge-transfer resonances near the VIS laser wavelength at the gold-molecule interface may enhance the MIR non-resonant susceptibility, in the same way that it has been shown to contribute to Raman scattering enhancement in SERS [55, 56, 57]. Proving this hypothesis will require tuning the VIS wavelength over a broad range, as in [58, 54], or controllably shifting the energy of the charge-transfer state by molecular engineering [59] or application of electric fields [60] – which can also break inversion symmetry and thereby increase second-order nonlinearities [61]. An SFG study of dodecanethiol molecules adsorbed on Au, Ag, and Pt surfaces [58] suggested that the Fermi level of the substrate plays a key role in enhancing the non-resonant SFG signal. This enhancement could also be modulated by electron-withdrawing groups on the molecules.
@@
All data presented above establish a reliable plasmonic platform combined with a simple laser technique to investigate the combined MIR-VIS nonlinear response of metallic nanogaps, which are taking central roles in a growing number of research and application fields, such as photo-catalysis [62], molecular and nano opto-electronics [63, 64], photodetection [65] and upconversion [66], nanoscale thermal transport [67, 68], polaritonics [69], vibrational strong coupling [70], 2D material opto-electronics [71], biosensors [72], chemical fingerprinting [73], single-molecule spectroscopy [74, 75], intra-molecular dynamics [76], etc. In Fig. S6, we show SFG spectra acquired in a single CCD exposure under fast continuous tuning of the MIR laser, a useful capability for applications requiring high throughput, such as hyperspectral far-field [40, 77, 78] or near-field [79, 80, 81] MIR and SFG imaging.
@@
The platform’s versatility is validated across multiple molecules (BPhT, 4-NTP, cn-BPhT), showing reproducible, chemically specific signals from nanoscale volumes. The absence of geometric phase-matching requirements, the use of a tunable cw MIR laser, and the operation under ambient conditions pave the way for chip-integrated molecular sensing and single-molecule nonlinear spectroscopy. Given the high signal-to-noise ratio currently obtained with a few hundred molecules in the probed volume (see Fig. 1e), the detection of SFG from a single molecule seems within reach [82], using, for example, host-guest chemistry [83] or DNA origami techniques [84, 85] to trap a controlled number of molecules. In this regime, the anharmonicity of the molecular vibrations has been predicted to result in non-classical anti-bunching of upconverted photons [86], opening new applications in quantum optics and quantum-enhanced spectroscopy, not accessible to the recently developed single-molecule IR spectroscopy methods that leverage incoherent (e.g., fluorescence) readout [87, 88, 89, 90, 91, 92, 93, 94].
@@ Acknowledgments
F. C. thanks Roberto A. Boto for insightful exchanges about the DFT modelling. Z. X. and C. G. thank Matthias Godejohann for valuable assistance with the MIR laser, and the EPFL Cmi and IPHYS staff for technical support. This research received funding from the Swiss National Science Foundation (SNSF project No. 214993) and the European Union’s Horizon 2020 research and innovation program under Grant Agreement No. 820196 (ERC CoG ‘QTONE’)

### v2

@@ Discussion
Several open questions remain. First, the SFG spectrum from BPhT presents significant deviations from that predicted by DFT calculations. Could this be related to the embedding of the molecules in a plasmonic nanogap? Is the discrepancy a signature of vacuum electromagnetic field modification [44], or rather linked to chemical interaction altering the geometry or electronic states of the molecules? Or is it related to other contributions not yet accounted for, such as from charge-transfer resonances? Unfortunately, we could not find broadband SFG spectra of BPhT from flat gold films in the literature to serve as a comparison. We attempted the DFT simulation of a more realistic geometry with molecules sandwiched between two gold planes using the model of Ref. [76], yet this did not yield better agreement with the data. Such complex metal-molecule interfaces are still challenging systems to accurately model with DFT.
@@
Second, the strength of the non-resonant varies markedly compared to the resonant vibrational response when changing the molecular species. This observation challenges the assumption that the non-resonant is an intrinsic property of the gold surface [77]. We speculate that the existence of charge-transfer resonances near the VIS laser wavelength at the gold-molecule interface may enhance the MIR non-resonant susceptibility, in the same way that it has been shown to contribute to Raman scattering enhancement in SERS [78, 79, 80]. Proving this hypothesis will require tuning the VIS wavelength over a broad range, as in [81, 77], or controllably shifting the energy of the charge-transfer state by molecular engineering [82] or application of electric fields [83] – which can also break inversion symmetry and thereby increase second-order nonlinearities [84]. An SFG study of dodecanethiol molecules adsorbed on Au, Ag, and Pt surfaces [81] suggested that the Fermi level of the substrate plays a key role in enhancing the non-resonant SFG signal. This enhancement could also be modulated by electron-withdrawing groups on the molecules.
Conclusion and perspectives
@@
All data presented above establish a reliable plasmonic platform combined with a simple laser technique to investigate the combined MIR-VIS nonlinear response of metallic nanogaps. Such platforms are taking central roles in a growing number of research and application fields, such as photo-catalysis [85], molecular and nano opto-electronics [86, 87], direct photodetection [88], frequency upconversion [89, 38, 90, 91], nanoscale thermal transport [92, 93], polaritonics [94], vibrational strong coupling [95], 2D material opto-electronics [96], biosensors [97], chemical fingerprinting [98], single-molecule spectroscopy [99, 100], intra-molecular dynamics [101], etc. In Fig. S6, we show SFG spectra acquired in a single CCD exposure under fast continuous tuning of the MIR laser, a useful capability for applications requiring high throughput, such as hyperspectral far-field [45, 102, 103] or near-field [104, 105, 106] MIR and SFG imaging.
@@
The platform’s versatility is validated across multiple molecules (BPhT, 4-NTP, cn-BPhT), showing reproducible and chemically specific signals from nanoscale volumes. Importantly, the absence of geometric phase-matching requirements, the use of a tunable cw MIR laser, and the operation under ambient conditions pave the way for chip-integrated molecular sensing.Given the high signal-to-noise ratio currently obtained with a few hundred molecules in the probed volume (see Fig. 1e), the detection of SFG from a single molecule seems within reach [107] (see Fig. S7). This could be realized using host-guest chemistry [108] or DNA origami techniques [109, 63] to trap a controlled number of molecules. In this regime, the anharmonicity of the molecular vibrations has been predicted to result in non-classical anti-bunching of upconverted photons [110]. Such quantum signatures would open new applications in quantum optics and quantum-enhanced spectroscopy, which are not accessible to recently developed single-molecule IR spectroscopy methods based on incoherent (e.g., fluorescence) readout [111, 112, 71, 113, 73, 72, 114, 115, 116].
@@ Acknowledgments
F. C. thanks Roberto A. Boto for insightful exchanges about the DFT modelling. Z. X. and C. G. thank Matthias Godejohann for valuable assistance with the MIR laser, the EPFL Cmi and IPHYS staff for technical support, and Arnaud Jollivet for earlier contributions to the setup. This research received funding from the Swiss National Science Foundation (SNSF project No. 214993) and the European Union’s Horizon 2020 research and innovation program under Grant Agreement No. 820196 (ERC CoG ‘QTONE’)